# Mapping Trustworthiness in Large Language Models: A Bibliometric Analysis Bridging Theory to Practice

### v1

@@ 4.8. The LLM trust/trustworthiness definitions proposed by different authors in the literature
Table 9. Frequency of Trustworthiness-Related Terms
4.9. Practical LLM trustworthiness-enhancing techniques found in the literature
From the 68 documents analysed, 19 unique documents contribute to the creation of the Table LABEL:tab:trust_enhancing, which provides strategies for enhancing trustworthiness in LLM. We found 20 different strategies, categorised by their LLM lifecycle phase – either pre-training, post-training or inference phase – the actor responsible for their implementation – either developer or user – and the trustworthiness attributes associated with them – e.g. transparency, explainability, reliability.
In the first place, with 6 different documents referring to this strategy, is (supervised) fine-tuning (SFT), in the post-training phase. This strategy is the responsibility of the LLM system developer and focuses mainly on reliability, robustness and fairness. This is justified by the boom of chatGPT, which is a fine-tuned version of the GPT-3 or GPT-4 pre-trained models. SFT has enabled models to become helpful assistants, that is, only the pre-trained models themselves were not the desired chatbot that chatGPT came to be. Nowadays, there is extensive documentation such as OpenAI fine-tuning docs, and platforms such as Google Colab, which have facilitated the availability of hardware and environment to allow fine-tuning by anyone anywhere in the world.
@@ 4.9. Practical LLM trustworthiness-enhancing techniques found in the literature
Developers have primary responsibility for trustworthiness (16/20 techniques), while users contribute minimally (4/20, e.g., human-in-the-loop feedback). This raises concerns about scalability, as techniques such as stakeholder feedback (Entry 5) require sustained user engagement, which is rarely seen in real-world deployments.
@@
Users can improve the trustworthiness of AI-LLM systems mainly in the inference phase, through human-in-the-loop (evaluating the output), stakeholder feedback (expert feedback), prompt engineering (implementing prompt techniques) and Reinforcement Learning from Human Feedback (RLHF) (providing feedback/labelling to the generated output).
Most techniques (12/20) target the post-training phase (e.g. fine-tuning, RAG, XAI methods). This overemphasis on post-training strategies rather than pre-training (e.g., bias mitigation through pre-training data curation) should not always suggest that developers are attempting to address reliability risks only after deployment. This is due to the fact that the pre-training phase (e.g., data gathering, training) typically requires millions of dollars and supercomputers. However, it is suggested from previous RQs that trust issues should be addressed before deployment, as early failures can severely damage user confidence and adoption, making it difficult for a platform to regain its image and user trust (Martell et al., 2024). In this study, we suggest that the model has passed the deployment phase only when it is in the inference phase.
Explainability and transparency dominate (14/20 entries), driven by techniques such as contextual explanations (Entry 8) and uncertainty expressions (Entry 11). In contrast, security (Entry 16: blockchain) and ethics (Entry 13: audits) receive limited attention, indicating a misalignment with regulatory priorities such as the EU AI Act’s emphasis on systemic risk management.
The strategies identified suggest the need for integrated frameworks that unify technical interventions (e.g. RAG, knowledge graphs) with governance mechanisms (e.g. post-release audits, human-in-the-loop) across the LLM lifecycle, with a strong adherence to legislations.
Efforts such as the EU AI Act From the policy making perspective, it is pressing to enact regulations to mandate transparency and accountability in the development and deployment of LLMs.
Table 10. Trustworthiness-Enhancing Strategies in LLMs
# Trustworthiness-Enhancing Strategy LLM Lifecycle Phase Responsibility Trustworthiness Attributes Citations
1 (Supervised) Fine-tuning Post-training Developer Reliability, Robustness, Fairness (Jha et al., 2023) (Kilhoffer and Bashir, 2024) (Barman et al., 2024) (Ressel et al., 2024) (Chen and Shu, 2024) (Roumeliotis et al., 2025)
2 RAG Post-training Developer Faithfulness, Transparency, Reliability (Jha et al., 2023) (Lan et al., 2024) (Ressel et al., 2024) (Hannah et al., 2025)
3 Training Pre-training Developer Robustness, Accuracy (Bhattacharya et al., 2023)
4 Human-in-the-loop Inference User Fairness, Transparency, User Trust (Wang et al., 2024b) (Chen and Shu, 2024)
5 Stakeholder feedback Inference User Trust Calibration, Social Acceptability (Wang et al., 2024b)
6 Prompt engineering Inference User Safety, Controllability, Fairness (Mousavi et al., 2024) (Hannah et al., 2025)
7 Evaluation Post-training Developer Accountability, Transparency (Mousavi et al., 2024) (Nag et al., 2023)
8 Contextual Explanation Pre-training/Post-training Developer Explainability, Transparency (Chari et al., 2023)
9 In-context Learning (ICL) Post-training Developer Adaptability, Efficiency, Fairness (Zhang et al., 2024) (Barman et al., 2024)
10 RAG-Ex: a model and language-agnostic explanation framework Post-training Developer Explainability, Transparency (Sudhi et al., 2024)
11 Uncertainty Expression in LLMs Post-training Developer Transparency, Explainability, Reliability (Kim, 2024)
12 LLMs’ self-explanations (i.e., explanations of their own answers and behaviors) Post-training Developer Explainability, Interpretability (Kim, 2024)
13 Post-release audit tool as an evaluative framework Post-training Developer Transparency, Accountability, Explainability, Ethics (Fernandez Nieto, 2024)
14 Reinforcement Learning from Human Feedback (RLHF) Inference Developer/User Alignment (Barman et al., 2024) (Chen and Shu, 2024)
15 Chain of Thought (CoT) Inference Developer Explainability (Barman et al., 2024) (Chen and Shu, 2024)
16 Blockchain Post-training Developer Security, Transparency, Tamper-resistance (Fan et al., 2024)
17 Knowledge Graph Post-training Developer Transparency, Explainability (Chen and Shu, 2024) (Hannah et al., 2025)
18 (Training) Data curation Pre-training Developer Fairness, Bias Mitigation (Chen and Shu, 2024)
19 (Access to external) Tools (e.g., access to search engines online) Inference Developer Explainability (Chen and Shu, 2024)
20 eXplainable AI (XAI) techniques – model simplification (LIME), perturbation-based methods (SHAP), gradient-based approaches (InputXGradient, Grad-CAM), Layer-wise Relevance Propagation (LRP) Post-training Developer Explainability, Interpretability, Transparency (Mersha et al., 2025)
5. Limitations
@@
While bibliometric analysis provides valuable insights into research trends, it has notable limitations. Its emphasis on publication and citation counts does not always capture the depth, quality or maturity of a field of research (Donthu et al., 2021). High citation rates can indicate popularity rather than true intellectual impact, and biases in citation patterns can distort impact (Donthu et al., 2021).
@@
To address these concerns, this study combines bibliometric techniques with qualitative assessments of articles with high focus on discussing trustworthiness (i.e., studies with ”trust*” in the abstract). By integrating both perspectives, we aim to provide a more comprehensive understanding of LLM trustworthiness research, including also more detailed, qualitative analysis of select papers in addition to the quantitative bibliometric approach.
@@
6. Final Remarks
This study provides a bibliometric analysis of trustworthiness in Large Language Models (LLMs), highlighting key research trends, influential papers and gaps in the literature. The findings reveal a shift from traditional AI ethics discussions towards trustworthiness frameworks, but a lack of consensus and practical implementation strategies persists. While various approaches such as transparency, accountability and governance mechanisms have been proposed, enforceable regulations and standardised frameworks remain necessary for real-world impact.
Future research should focus on bridging the gap between theoretical taxonomies and actionable trustworthiness-enhancing techniques to ensure their practical application in LLM development and deployment. In addition, interdisciplinary collaboration is essential to reconcile ethical principles with technical and regulatory considerations. Given the rapid evolution of LLMs and their increasing societal impact, ongoing efforts must prioritise both responsible innovation and policy-driven accountability. Ultimately, achieving trustworthy AI will require not only technological advances, but also robust governance, public engagement, and industry-wide commitment to ethical AI development.
Acknowledgements.

### v2

@@ 4 Wang C A Survey on Evaluation of Large Language Models [24]
From the 68 documents analysed, 19 unique documents contribute to the creation of the Table V, which provides strategies for enhancing trustworthiness in LLM. We found 20 different strategies, categorised by their LLM lifecycle phase – either pre-training, post-training or inference phase – the actor responsible for their implementation – either developer or user – and the trustworthiness attributes associated with them – e.g. transparency, explainability, reliability.
TABLE V: Trustworthiness-Enhancing Strategies in LLMs
# Strategy LLM Lifecycle Phase Responsibility Trustworthiness Attributes Citations
1 (Supervised) Fine-tuning Post-training Developer Reliability, Robustness, Fairness [26, 70, 71, 52, 72, 73]
2 RAG Post-training Developer Faithfulness, Transparency, Reliability [26, 74, 52, 75]
3 Training Pre-training Developer Robustness, Accuracy [76]
4 Human-in-the-loop Inference User Fairness, Transparency, User Trust [77, 72]
5 Stakeholder feedback Inference User Trust Calibration, Social Acceptability [77]
6 Prompt engineering Inference User Safety, Controllability, Fairness [78, 75]
7 Evaluation Post-training Developer Accountability, Transparency [78, 79]
8 Contextual Explanation Pre-/Post-training Developer Explainability, Transparency [80]
9 In-context Learning (ICL) Post-training Developer Adaptability, Efficiency, Fairness [81, 71]
10 RAG-Ex Framework Post-training Developer Explainability, Transparency [82]
11 Uncertainty Expression Post-training Developer Transparency, Explainability, Reliability [83]
12 Self-explanations Post-training Developer Explainability, Interpretability [83]
13 Post-release audits Post-training Developer Transparency, Accountability, Explainability, Ethics [84]
14 RLHF Inference Developer/User Alignment [71, 72]
15 Chain of Thought (CoT) Inference Developer Explainability [71, 72]
16 Blockchain Post-training Developer Security, Transparency, Tamper-resistance [85]
17 Knowledge Graph Post-training Developer Transparency, Explainability [72, 75]
18 Data curation Pre-training Developer Fairness, Bias Mitigation [72]
19 External Tools (e.g., search engines) Inference Developer Explainability [72]
20 XAI techniques (LIME, SHAP, etc.) Post-training Developer Explainability, Interpretability, Transparency [86]
In the first place, with 6 different documents referring to this strategy, is (supervised) fine-tuning (SFT), in the post-training phase. This strategy is the responsibility of the LLM system developer and focuses mainly on reliability, robustness and fairness. SFT has enabled models to become helpful assistants, that is, only the pre-trained models themselves were not the desired chatbot that ChatGPT came to be.
@@ 18 Data curation Pre-training Developer Fairness, Bias Mitigation [72]
Developers have primary responsibility for trustworthiness (16/20 techniques), while users contribute minimally (4/20, e.g., human-in-the-loop feedback). This raises concerns about responsibility balance, as the liability of positive trustworthiness outcomes relies mostly in the developers hands. Users can improve the trustworthiness of AI-LLM systems mainly in the inference phase, through human-in-the-loop (evaluating the output), stakeholder feedback (expert feedback), prompt engineering (implementing prompt techniques) and Reinforcement Learning from Human Feedback (RLHF) (providing feedback/labelling to the generated output).
Most techniques (12/20) target the post-training phase (e.g. fine-tuning, RAG, XAI methods). Regarding trustworthiness attributes, explainability and transparency come out on top (14/20 entries), driven by techniques such as uncertainty expressions (Entry 11) and knowledge graph (Entry 17). In contrast, security (Entry 16: blockchain) and ethics (Entry 13: audits) receive limited attention.
V Limitations
@@
While bibliometric analysis provides valuable insights into research trends, it has notable limitations. Its emphasis on publication and citation counts does not always capture the depth, quality or maturity of a field of research [16]. High citation rates can indicate popularity rather than true intellectual impact, and biases in citation patterns can distort impact [16].
@@
To address these concerns, this study combines bibliometric techniques with qualitative assessments of articles with high focus on discussing trustworthiness (i.e., studies with “trust*” in the abstract). By integrating both perspectives, we aim to provide a more comprehensive understanding of LLM trustworthiness research, including also more detailed, qualitative analysis of select papers in addition to the quantitative bibliometric approach.
@@
VI Discussion and Conclusions
As shown in Figure 1, the sharp increase in publications after 2022 can be attributed to advances such as GPT-3/4, regulatory debates (e.g., EU AI Act), and societal concerns about LLM trustworthiness and ethical outcomes. This highlights the increase interest researchers have in trust/trustworthiness and ethical LLM and genAI research. This perceived trend is expected to continue, reflecting the growing attractiveness of the field to both established researchers and newcomers.
Figure 2 shows articles with high global citations, indicating a significant impact and influence across disciplines [16]. Kasneci et al. [19] explore the opportunities and challenges of using LLMs like ChatGPT in education, highlighting their potential to personalize learning, create educational content, and support both students and teachers. However, it also addresses critical risks such as bias, over-reliance, ethical concerns, and sustainability, emphasizing the need for responsible integration, teacher training, and ongoing research. The authors advocate for a balanced approach that uses LLMs to improve education while providing ethical use, transparency, and equitable access. Ongoing efforts to perceive the use of LLMs in education highlights how academia is concerned about this disruptive tool in educational settings.
Dwivedi et al. [20] explores the multidisciplinary impact of ChatGPT, highlighting its potential to enhance productivity across industries while raising concerns about ethics, misinformation, and transparency. It discusses its role in research, business, and education, emphasizing both opportunities and challenges, including AI’s effect on academic integrity and hybrid work. The authors propose multiple future research directions, e.g., on AI governance, human-AI collaboration, and responsible implementation.
Sallam Malik [21] systematically reviews the potential applications and limitations of ChatGPT in healthcare. It highlights the benefits of ChatGPT in scientific writing, research efficiency, medical education, and clinical workflows, while addressing concerns such as misinformation, bias, ethical and legal challenges, and safety risks. The study emphasises the need for responsible use of AI in healthcare and calls for regulatory guidelines and stakeholder collaboration to ensure ethical and effective implementation.
The three highly cited documents analysed present investigations of the specific adoption and impact of ChatGPT in different domains – education [19], research, business, and education [20], and healthcare [21]. This shows how the research community has responded to the rapid emergence of ChatGPT in several domains. Furthermore, all three top documents have in common the need to ensure the ethical use, transparency and responsible implementation of AI.
The results shown in Figure 3 indicate a progressive shift in the discourse from general AI ethics toward more specialized concerns about LLM trustworthiness. The focus on ‘AI ethics’ (2023-2024) has shifted to concerns about ‘disinformation’ (2025-2025), reflecting the fact that LLMs are rapidly replacing traditional search engines and are primarily used as chatbots, particularly in question-answering (QA) settings. As users increasingly rely on LLMs rather than traditional search engines to retrieve information, the risk of misinformation and manipulated content has become a central issue. This typical type of use of LLMs suggests why discussions of AI ethics have evolved to disinformation.
The term ‘trustworthiness’ (2025-2025) stems from the term ‘responsible AI’ (2023-2024), suggesting the discourse shift from traditional responsible AI principles to trustworthiness in genAI. This suggests that responsible AI discussions have been rebranded and adapted to meet the LLM challenges. Moreover, AI ethics research focus has shifted from abstract guiding principles – e.g., [11, 2, 1, 12]. – to abstract LLM trustworthiness taxonomies – e.g., [6, 8, 7].
Researchers unfamiliar with ethical aspects in AI are being attracted to explore it but in the LLM and genAI hype. Thus, they are coming up with terminologies and theoretical approaches, that can be seen as yet another set of AI ethics guidelines. This abrupt interest is possibly hindering the actual advancements towards more ethical AI, as it is masking the debate towards a “new” set of theoretical guidelines, but not providing practical guidance for the developers, the ones ultimately responsible for the development of ethical AI or trustworthy LLM [87]. In other words, this perceived rebranding of responsible AI and AI ethics to trustworthiness and disinformation is for the most part failing in proposing useful unified terminologies, frameworks, or practical guidance for developers to operationalise principles.
There is a growing need for enforceable guidelines beyond non binding ethical commitments. Even more evident after big tech companies such as Google withdrawing its AI ethical principles in mass surveillance and AI weapons [88]. Consequently, there is an urge for actual hard laws that can govern the development and use of these systems, that can move beyond the EU AI Act [89], due to the fact that it only propose risk level assessments, and links to The European Commission’s High-level expert group on AI (AI HLEG) [90]. The later has a set of trustworthy AI principles, however, it is non binding. Despite the fact that a binding document links to a non binding one, does not automatically transform the AI HLEG into hard law. On top of that, an unified set of practical guidance are still lacking, after several years of debate – at least since 2019 [87].
Despite this progress, fragmentation remains, with multiple authors competing with definitions of LLM trustworthiness rather than a single, unified framework. The continued proliferation of non-binding framework proposals and terminologies by singular researchers highlights the urgency of bridging the gap between theoretical AI ethics and its real-world implementation through unified frameworks and legislation. This observation goes in line with tech ethics limitations identified by Green’s [91] where a vagueness and focus on individuals leads to ethics washing, stakeholders ease public voice and resist to governmental legislation by embracing the ethical discourse, without real commitment.
Figure 4 shows the most frequent keywords that appeared in the analysed literature on LLMs, genAI, trust and ethics. ‘ChatGPT’ is the most recurrent term as it dominates as the most popular LLM in the literature, while ‘health’ and ‘education’ are the domains most explored. Furthermore, ‘privacy’, ‘ethics’, ‘bias’ are the most relevant regarding ethical issues. This indicates that most of the studies are investigating how privacy and bias are the most explored terms in health and education through the use of ChatGPT.
From the coupling map, the studies identified in Figure 5 and Table I highlight the applications of LLMs across multiple domains and the ethical, practical, and evaluative challenges. Martinez-Maldonado and colleagues [22] systematically review the use of LLMs in education and identify nine key applications, such as grading, content generation and feedback, while cautioning about replicability, transparency and ethical concerns. Park and co-authors [23] investigate the impact of misinformation generated by ChatGPT on tourism decision-making, demonstrating how trust in AI can be easily manipulated when incorrect information is presented. Similarly, Sallam [21] examines AI in healthcare, pinpointing its potential to improve medical research, education and clinical workflows, as well as its potential risks such as bias, hallucination, plagiarism and cybersecurity threats.
Wang et al. [24] provide a meta-analysis of AI evaluation methods and argue for a more structured approach to assessing their performance, ethical considerations and real-world impact in different domains. Herrera et al. [25] explore XAI, pointing the urgent need for transparency in AI decision-making and proposes 28 open problems for achieving better interpretability and accountability in AI systems.
Overall, these authors are exploring trust in domains involving education, tourism, healthcare, stressing how LLMs are reshaping a multitude of sectors in everyday life. Furthermore, the studies emphasize how evaluation methods and transparency are closely related to trust.
In an attempt to find LLM trustworthiness definitions from the literature, we devised Table II, from which we point that several authors are building on Mayer et al.’s [10] organizational trust model, indicating a deficiency in the development of a new framework specifically made for AI and LLM systems. While Mayer’s model is broadly adopted in academia, it was devised before the widespread adoption of AI technologies (1995). Thus, adapting a 30-year-old framework to the AI era raises important questions. Similarly to AI ethics, where soft law recommendations did not pose serious changes, it is worth discussing whether trustworthiness in LLMs should instead be governed by hard laws that unify terms and recommendations.
The concept of trust calibration – especially the risk of over-trust and under-trust described by Martell et al. [59] – reinforces the importance of addressing trustworthiness early in the deployment process. As the case of Microsoft’s Tay Bot demonstrates, early missteps can permanently damage user trust, making recovery almost impossible. The findings also highlight that users are inherently vulnerable when interacting with LLMs and generative AI systems, as they rely on these systems with the expectation of goal achievement despite possible misinformation or bias.
Furthermore, the more familiar we are with a tool, the more we tend to trust them [53]. Other authors state that trust in AI-LLM can be translated as the degree to which users and institutions are confident that the data used by, and the information generated by, AI-LLM is impartial, i.e. unbiased [54].
Two more trust related dimensions are presented in Table II: cognitive trust (rational evaluation of a system’s reliability and accuracy) and affective trust (emotional confidence in its benevolence and alignment with user interests) [66]. Cognitive trust, anchored in perceived credibility, facilitates affective trust through sustained reliance, as users transfer trust from the platform’s technical competence to its perceived ethical intent [66, 67]. This interdependence highlights how platform reliability shapes users’ perceptions of AI-generated content, with technical accuracy (‘reliable’) and empathetic design (‘credible’) jointly calibrating holistic trust.
In [54], the authors state three principles of AI-LLM: fair, unbiased, and ethical. On top of that, they relate these principles with the work done within the European Commission on AI “Building Trust in Human-Centric Artificial Intelligence” [34], which states that trustworthy AI should not only be lawful, ethical, and robust, but also give humans the ability to have control over its use, draw from and provide reliable data, offer security and privacy to its users, and be transparent, accountable, and inclusive. The efforts of the European Commision on AI is better translated in the EU AI Act [89, 92], where a risk level assessment for AI systems is provided. We briefly describe the EU AI Act’s risk level in Table VI.
TABLE VI: Risk Levels in the EU AI Act
Risk Level Description
Unacceptable Risk AI systems that are prohibited due to threats to fundamental rights, e.g., social scoring, manipulation.
High Risk AI systems that impact safety or fundamental rights, e.g., biometric ID, credit scoring, law enforcement.
Limited Risk AI systems with transparency requirements, e.g., chatbots, AI-generated content.
Minimal Risk AI systems with no specific regulations, e.g., spam filters, AI-driven recommendations.
In addition, within the EU AI Act there is reference to a non-binding document, also provided by the European Commission, that aims to provide a set of trustworthy principles in AI. This non-binding document is the 2019 Ethics Guidelines for Trustworthy AI developed by the AI High-Level Expert Group (AI HLEG) [90]. We shortly present its principles in Table VII. This Table reinforces terms already discussed in Table IV, showing a level of alignment between academic and policy perspectives.
TABLE VII: AI HLEG Principles for Trustworthy AI – 2019
Principle Description
Finally, in [69], there is a critic towards the use of the terms trustworthy and ethical AI, as they can be used to humanize, i.e., anthropomorphise AI. This critic is valid insofar every Information System artefact, such as LLM and AI systems, are devised by humans, and they are programmatically designed to fulfil a set of subjective or objective desires implied by the developers or organizations they are inserted.
In Table IV Transparency and Explainability appear as the first and second most frequently mentioned characteristics, respectively, highlighting how they are enabling principles, i.e., without transparency and explainability it is impossible for other trustworthiness or ethical principles to exist [93]. Reliability is related to Ability and Predicability in the ABI+ framework [10]. Furthermore, Ethics, Privacy, and Fairness highlight concerns related to Integrity, i.e., adherence to ethical standards.
The techniques most frequently mentioned in the literature, as shown in Table V are SFT and RAG, which stand out for their technical feasibility. Furthermore, most of the strategies identified are post-training. This overemphasis on post-training strategies, rather than addressing issues earlier in the pre-training phase (e.g., bias mitigation through pre-training data curation), should not necessarily be interpreted as a lack of concern for preventive risk mitigation. On the contrary, it reflects the resource-intensive nature of the pre-training phase, including data collection and model training. In addition, the developer is responsible for most of the strategies, highlighting the crucial role of practitioners in operationalising trustworthy LLM and the need to provide them with practical guidance.
Regarding the distribution of trustworthiness attributes identified, it indicates a misalignment with regulatory priorities such as the risk levels of the EU AI Act, e.g., security appearing only once. Further, the strategies identified suggest the need for integrated frameworks that unify technical interventions (e.g. RAG, knowledge graphs) with governance mechanisms (e.g. post-release audits, human-in-the-loop) across the LLM lifecycle, with a strong adherence to legislations. From a policy perspective, efforts such as the EU AI Act are steps in the right direction, but there remains an urgent need for stronger regulations that mandate transparency and accountability in the development and deployment of LLMs.
In conclusion, despite the growing body of literature on LLM trustworthiness, there is still no unified set of practical guidelines to help developers of LLM-based systems mitigate risks throughout the development and deployment process. To bridge this gap, we conducted a bibliometric analysis and manual reading of papers related to trustworthiness in LLMs, highlighting research trends, definitions and practical guidance. The findings reveal a shift from traditional AI ethics discussions towards trustworthiness frameworks, but a lack of consensus and practical implementation strategies persists. Most common terms related to trustworthiness extracted from reviewed papers are transparency, explainability, reliability, ethics and privacy. While various technical strategies were idenified, enforceable regulations and standardised frameworks remain necessary for real-world impact. Multiple authors are competing for LLM trustworthiness terminologies rather than unified frameworks and legislation, this fragmentation and vagueness leads to risks of “ethics washing,” where stakeholders ease public voice and resist to legislation by embracing the ethical discourse, without real commitment.
Future work includes implementing and evaluating individual trustworthiness-enhancing strategies in LLMs identified. Moreover, different combinations of technical strategies, such as RAG, knowledge graph, XAI and fine-tuning, would shed some light on the trustworthiness practices. Furthermore, adequate training and practical guidance for developers in the operationalisation of trustworthiness attributes should be provided. In addition, interdisciplinary collaboration is essential to reconcile ethical principles with technical and regulatory considerations, thus involving potential stakeholders as well as users of the system. Ultimately, progress towards trustworthy AI will require not only technological implementations, but also unified legislation, practical frameworks, public and industry engagement to ethical AI development.
Acknowledgment